# FireFly-S: Exploiting Dual-Side Sparsity for Spiking Neural Networks Acceleration with Reconfigurable Spatial Architecture

### v1

@@ VI-B Parallelism Setup
As outlined in Section V, FireFly-S leverages parallelism to enhance throughput and computational efficiency. Determining the optimal parallelism levels for the inter-layer pipelined architecture is crucial to maximizing performance within device constraints. This optimization is achieved through a two-step process.
@@ VI-C Resource and Performance Evaluation
Table V offers a detailed breakdown of resource usage by each module within a single layer of FireFly-S, indicating that the Dataflow Orchestrator and Dual-side Sparsity Detector are the primary consumers of resources when the Maxpooling module is inactive. Note that the reported resource usage for the Dual-side Sparsity Detector, except for BRAMs, is calculated by dividing by the parallelism level (26 in this experiment) Therefore, while increasing parallelism in FireFly-S can enhance performance, it will also raise the consumption of LUTs and FFs, potentially reaching the resource limits.
@@
Furthermore, Table VI highlights prevalent data precisions in SNN implementations, specifically 16-bit and 8-bit. While the study by Gerlinghoff et al. [25] applies 3-bit quantization to LeNet5, it utilizes 6-bit precision for more complex models like AlexNet and VGG11 to maintain acceptable accuracy levels. In a similar vein, Panchapakesan et al. [8] implement percentile-based quantization in SyncNN, achieving 4-bit precision for LeNet-L and 8-bit for NIN and VGG13. However, attempts to apply 4-bit quantization to NiN and VGG13 resulted in significant accuracy reductions of 7.86% and 8.96%, respectively. In contrast, the FireFly-S model maintains high accuracy not only with 4-bit quantization across various models but also achieves notable sparsity levels of 93.82% in SCNN5, 93.26% in SCNN7, and 86.12% in SCNN9. Additionally, our hardware architecture uniquely supports dual-side sparsity exploitation, distinguishing it from previous implementations.
@@ VII Conclusions
In this work, we introduce a high-energy-efficient and reconfigurable hardware accelerator for SNNs, leveraging a software-hardware co-design approach to optimize energy efficiency during SNN inference. On the software side, our approach includes a quantization-aware and pruning-aware training method, which enables the training of SNN models at 4-bit quantization and 85-95% sparsity with minimal accuracy loss. This training strategy not only allows the deployment of large SNN models on-chip but also enhances calculation logic utilization by introducing weight sparsity through pruning. On the hardware side, we capitalize on the natural sparsity of spikes inherent to SNNs and the engineered sparsity of weights to accelerate processing. This is effectively exploited by our Dual-side Sparsity Detector, which accelerates processing by leveraging both types of sparsity and skipping the unnecessary calculations. Additionally, our architecture is designed to be fully pipelined, enhancing throughput. The Dataflow Orchestrator plays a critical role in maintaining consistent dataflow between layers, minimizing stalls, and simplifying logic to reduce resource consumption. We validate the reconfigurability of our FireFly-S accelerator by evaluating various deep SNN models across multiple datasets. Utilizing commercially available FPGA edge devices offers a more practical and feasible solution compared to other existing hardware platforms. Future work will focus on enhancing the SNN hardware infrastructure further, aiming for ultra-high compression rates and addressing the decline in efficiency that occurs when sparsity levels are excessively high in Bitmap representation.

### v2

@@ VI-B Parallelism Setup
As outlined in Section V, FireFly-S leverages parallelism to enhance throughput and computational efficiency. Optimizing the parallelism levels for the inter-layer pipelined architecture is crucial to maximizing performance within device constraints, which is achieved through a two-step process.
@@ VI-C Resource and Performance Evaluation
Table V offers a detailed breakdown of resource usage by each module within a single layer of FireFly-S, indicating that the Dataflow Orchestrator and Dual-side Sparsity Detector are the primary consumers of resources when the Maxpooling module is inactive. Note that the reported resource usage for the Dual-side Sparsity Detector, except for BRAMs, is calculated by dividing by the parallelism level (26 in this experiment). Therefore, while increasing parallelism in FireFly-S can enhance performance, it will also raise the consumption of LUTs and FFs, potentially reaching the resource limits.
@@
Furthermore, Table VI highlights prevalent data precisions in SNN implementations, specifically 16-bit and 8-bit. While the study by Gerlinghoff et al. [34] applies 3-bit quantization to LeNet5, it utilizes 6-bit precision for more complex models like AlexNet and VGG11 to maintain acceptable accuracy levels. In a similar vein, Panchapakesan et al. [8] implement percentile-based quantization in SyncNN, achieving 4-bit precision for LeNet-L and 8-bit for NIN and VGG13.However, applying 4-bit quantization to NIN and VGG13 resulted in significant accuracy reductions. As shown in Table VII, while SyncNN maintains relatively high accuracy with 4-bit quantization for LeNet-L, more complex models like VGG13 and NIN experience substantial accuracy drops, from 95.65% to 87.87% and from 88.09% to 80.33%, respectively.
In contrast, FireFly-S tends to maintain accuracy after applying 4-bit quantization, especially in more complex datasets like CIFAR10, where the accuracy drop is minimized (from 87.86% to 87.15% for SCNN9). These enhanced accuracy levels are primarily due to our quantization-aware and pruning-aware training strategy, which leverages learnable per-channel scales and a hardware-friendly, computation-error-aware approach. In some cases, it may even result in slight accuracy improvements, as seen with SCNN7, where accuracy increases from 87.12% to 87.87%. In addition to experiencing less accuracy drop compared to SyncNN with our quantization method, we also combine it with pruning to further compress the network. Specifically, we achieve sparsity levels of 93.82% for SCNN5, 93.26% for SCNN7, and 86.12% for SCNN9, while maintaining competitive accuracy loss. For example, SCNN9 experiences only a slight accuracy drop from 87.86% to 87.00% after both quantization and pruning. Notably, in the DVS-Gesture dataset, SCNN7 even shows an accuracy improvement, rising from 87.12% (without compression) to 92.05% after applying our compression techniques. These results highlight that our method not only offers a hardware-friendly approach but also achieves highly competitive compression rates while preserving accuracy, demonstrating our joint quantization and pruning strategy is well-suited for low-precision and hardware-optimized inference.
[b] SyncNN LeNet-L VGG13 NIN Acc.@MNIST Acc.@SVHN Acc.@CIFAR10 Base1 99.602 95.65 88.09 Quant 99.60 87.872 80.33 Ours SCNN5 SCNN7 SCNN9 Acc.@MNIST Acc.@DVSG Acc.@CIFAR10 Base 99.57 87.12 87.86 Quant 99.54 87.87 87.15 Prune ${\text{98.56}}_{\text{@93.19}}$3 ${\text{92.42}}_{\text{@92.62}}$ ${\text{86.57}}_{\text{@85.74}}$ Quant + Prune ${\text{98.03}}_{\text{@93.82}}$ ${\text{92.05}}_{\text{@93.26}}$ ${\text{87.00}}_{\text{@86.12}}$
TABLE VII: Compression and Accuracy Results Compared to SyncNN
  • 1
    The base accuracy reported for SyncNN refers to 8-bit quantization, not an uncompressed model.
  • 2
    Values are estimates based on the Fig. 9 of SyncNN[8], as no exact values were provided in the original text.
  • 3
    The symbol ”$\text{X}_{\text{@Y}}$” indicates that accuracy = Xand sparsity = Y%.
@@ VII Conclusions
In this work, we introduce a high-energy-efficient and reconfigurable hardware accelerator for SNNs, leveraging a software-hardware co-design approach to optimize energy efficiency during SNN inference. On the software side, our approach includes a quantization-aware and pruning-aware training method, which enables the training of SNN models at 4-bit quantization and 85-95% sparsity with competitive accuracy loss. This training strategy not only allows the deployment of large SNN models on-chip but also enhances calculation logic utilization by introducing weight sparsity through pruning. On the hardware side, we capitalize on the natural sparsity of spikes inherent to SNNs and the induced sparsity of pruned weights to accelerate processing. This is effectively exploited by our Dual-side Sparsity Detector, which accelerates processing by leveraging both types of sparsity and skipping the unnecessary calculations. Additionally, our architecture is designed to be fully pipelined, enhancing throughput. The Dataflow Orchestrator plays a critical role in maintaining consistent dataflow between layers, minimizing stalls, and simplifying logic to reduce resource consumption. We validate the reconfigurability of our FireFly-S accelerator by evaluating various deep SNN models across multiple datasets. Utilizing commercially available FPGA edge devices offers a more practical and feasible solution compared to other existing hardware platforms. Future work will focus on enhancing the SNN hardware infrastructure further, aiming for ultra-high compression rates and addressing efficiency declines at extreme sparsity levels in Bitmap representation.